# DRIFT: Deep Reinforcement Learning for Intelligent Floating Platforms Trajectories

### v1

@@ III-F Laboratory Experiment Setup
For real-world validation, we conducted experiments using the physical air bearings platform [22] located in the ZeroG Laboratory at the University of Luxembourg. This platform, with a weight of 5.32 Kg, radius and heigtht of 31 and 45 cm, is equipped with a Raspberry Pi 4 for onboard control and interfacing. The ZeroG Lab contains an Optitrack Motion Capture System (MCS) that precisely tracks the platform’s pose at a high frequency of 200 Hz. From these measurements, we derive linear and angular velocities, which form the core components of the state representation required for policy inference. Our experimental setup maintains a connection between a laptop, the MCS, and the FP through a local network. The laptop serves as the ROS (Robot Operating System) master node of the network, subscribing to the Optitrack node to acquire pose data and publishing the actions of the trained agents at a rate of 5 Hz. This action frequency is deliberately constrained to prevent damage to the solenoid valves controlling the thrusters on the floating platform.
Figure 1 illustrates the key components interacting during the simulated training and validation phase (on the left), and those interacting during the closed-loop control tests of the real FP system in the Lab (on the right).
@@ IV Experimental Setup
Our experiments encompass both simulation-based evaluations and real-world experimental validations.
IV-A Simulation-Based Evaluations
For the evaluation, each trained policy was tested across a diverse set of scenarios defined by various environmental conditions. The system and environmental disturbances are:
  • Action Noise (AN): a random disturbance force of $\pm$ $x$ N applied to every thruster;
  • Velocity Noise (VN): $\pm$ $x$ m/s added to the state velocities;
  • Uneven Floor (UF): $x$ N of force, added to simulate the floor unevenness, applied to the FP body throughout the episode, either with a constant direction or through a sinusoidal generated direction;
  • Torque Disturbance (TD): $x$ Nm of torque applied to the body’s center of mass;
  • Random Thrusters Failure (RTF): a zeroing mask over the output actions to simulate one or multiple thruster failures which remains the same throughout the episode.
IV-B Tasks and Episodes
We define the different task as:
  • Go to position: Starting from a random initial position in the plane, reach the given (x, y) coordinates;
  • Go to pose: Starting from a random initial position in the plane, reach the given pose (position and orientation $\theta$);
  • Track velocity: Track the given velocity vector to follow a trajectory.
IV-C Performance Metrics
@@
IV-D Real-World Experimental Validations
To validate the transferability of the simulation-trained policies to the physical FP, we conducted real-world experiments using the laboratory setup described in section III-F. Five test runs were conducted for each policy, starting from different initial conditions. The variability in initial conditions includes random position and orientation. The same simulation task-specific metrics, including achieved accuracy, energy efficiency and trajectory smoothness, were employed.
@@ V RESULTS
AN VN UF TD RTF PT5 PT2 PT1 OT5 OT2 OT1 ALV AAV AAS
(%) (m/s) (N) (N/m) (-) (%) (%) (%) (%) (%) (%) (m/s) (rad/s) (-)
Ideal RL - - - - - 64 34 6 94 89 73 0.08 0.12 0.29
LQR - - - - - 73 41 17 27 11 5 0.07 0.16 0.10
Velocity Noise RL - 0.02 - - - 64 30 7 94 90 72 0.08 0.12 0.31
RL - 0.04 - - - 61 21 6 94 89 66 0.09 0.13 0.31
LQR - 0.02 - - - 53 21 6 4 1 0 0.09 0.49 0.23
LQR - 0.04 - - - 14 3 0 2 1 0 0.15 0.56 0.29
Constant Torque RL - - - 0.05 - 63 24 2 94 86 61 0.08 0.12 0.35
LQR - - - 0.05 - 57 20 6 3 1 0 0.07 0.43 0.35
Constant Force RL - - 0.20 - - 63 29 7 94 90 74 0.09 0.12 0.30
RL - - 0.40 - - 52 19 5 94 89 72 0.09 0.12 0.31
LQR - - 0.20 - - 66 17 4 28 12 6 0.07 0.15 0.12
LQR - - 0.40 - - 23 0 0 30 13 6 0.08 0.16 0.15
Constant Force & Torque RL - - 0.20 0.05 - 62 24 5 94 86 61 0.08 0.12 0.35
LQR - - 0.20 0.05 - 13 2 0 3 1 0 0.07 0.44 0.32
Thruster Failures RL - - - - 1 32 15 6 70 55 36 0.10 0.12 0.28
RL - - - - 2 15 6 2 45 31 20 0.16 0.15 0.25
LQR - - - - 1 40 17 5 20 8 4 0.10 0.21 0.16
LQR - - - - 2 12 4 1 11 4 2 0.14 0.28 0.22
V-A Benchmark RL & LQR
In this section, we explore the behavior of an RL agent trained to perform the “go to pose” task, and compare it to the LQR controller. To characterize the controllers’ behaviors we expose them to a range of disturbances. Neither the RL agents nor the LQR are specifically adapted to incorporate methods from robust RL or robust optimal control theory. However, it should be noted that the RL agent was trained with some domain randomization to learn how to deal with force disturbances up to 0.25N. Both of them are evaluated in MuJoCo, with similarly randomized initial conditions. In Table II, each line corresponds to an experiment, with different disturbances being applied, and was acquired using 256 trajectories of 250 steps.
@@ V-A Benchmark RL & LQR
When considering the Velocity Noise (VN), it is observed that with the lowest noise level, the RL performances remain unchanged, while the LQR is struggling, in particular with the attitude control. With 0.04 m/s of noise, the performance of both controllers decreases. Yet, the RL controller is more resilient than the LQR controller to this kind of disturbance, even though it was not trained for it. In the interest of brevity, we do not report action noise value in the table, as we found their effect to be negligible on both controllers.
Furthermore, when examining the Torque Disturbance (TD) of 0.05N/m, equivalent to 1/6-th of the total torque capacity of the platform, the performance of both controllers experiences a noticeable reduction; particularly, for the LQR controller. A similar pattern is observed with the force disturbance (UF), which would be equivalent to an uneven floor in the lab. In this case, we start by applying 0.2N of force on the platform, equivalent to 1/5-th of its maximum thrust. In this case, the performance of both controllers is close to the ideal conditions, with a small performance drop of the LQR in fine positioning. When doubling it (0.4N), the RL policy remains close to its baseline, but the LQR performance decreases with it being unable to maintain positions under the 2.5cm threshold. Similar behaviors are observed upon the addition of both force and torque disturbances.
@@
Overall, while the LQR controller demonstrates greater efficiency and precision in position control with our current tuning, it encounters challenges when subjected to the selected range of disturbances. In contrast, RL exhibits a lower degree of energy conservation but offers stronger resilience when subject to a wide range of disturbances. It is possible that with a different cost function, better tuning of its weights, and a robust optimal control approach, the LQR becomes adept with these disturbances. Yet, the RL agent is not using a robust RL approach either, and domain randomization was only applied on force disturbances up to 0.25N which is less than the disturbances it can overcome.
@@ V-B ZeroG Laboratory
Figure 2: Comparison of the RL and LQR controller on two different initial poses in the ZeroG lab. Init 1, denotes the first initial pose, Init 2 the second initial pose. For the trajectories, the scale of the y-axis is represented as a log value for better visualization.
@@
For experiments with the real FP system, we report two “go to pose” trajectories for both the RL and LQR methods. The controllers are run on the FP which is connected to a constant air supply through a tether. This tether applies some light unknown disturbances such as a small torque and force to the platform. Moreover, the system velocities are derived from the optitrack system. The observed velocities include a minor noise and experience small delays due to network communication.
@@
Figure 2 illustrates the performance of each controller. The first row shows the trajectories of the FP, the second row the distance to the goal in position and in radiant. The first two columns have the rough same initial pose: Init1, while the two last share the same initial pose: Init2.
@@
From the last row, it is evident that the LQR controller converges faster in position than the RL controller. This aligns well with the behaviors observed in the simulation benchmark, with an LQR controller converging faster. However, it is also apparent that the LQR solution exhibits a minor overshoot. Such an observation is also inline with the simulation benchmark, as the uneven floor in the lab likely disrupts the LQR controller by applying a subtle constant force, preventing it from reaching its simulation baseline performance. Looking at the top row, we can see that the LQR is also overshooting a bit. Of course, the behavior can be adjusted by modifying the weights associated with the importance of the error in position in the cost matrix. It is also worth noting that the LQR controller is sensitive to the weights; smaller weights do not incentivize the FP motion toward the goal. In comparison to the simulation, it was deemed necessary to alter the weights of the LQR controller to yield a more aggressive approach to achieve satisfying performances. As for the RL agent, it is noticeable that the FP initially aligns its heading and then gradually converges toward the goal. Consistently with the results from the simulation, the RL controller is significantly more accurate in terms of heading while achieving a position accuracy similar to that of the LQR controller. Overall, both controllers performed well in the lab, reaching their expected performances.
@@ VI CONCLUSIONS
This study presents a comprehensive methodology for controlling a FP in 2D using deep reinforcement learning. The enhancements to our simulator extend its applicability, enabling the training of agents for complex tasks. The demonstrated effectiveness of the PPO algorithm, with seamless transfer behavior from the simulation to the FP system, highlights its potential for autonomous navigation in space. Furthermore, our comparison with traditional optimal control algorithms underscores the potential of the proposed approach in terms of absolute performance and especially when unknown stochastic events alter the environmental conditions.
In the future, we will experiment with more complex networks such as LSTMs or Transformers to see if they can cope better with the disturbances and in particular RTFs. We will also extend this framework to train agents to walk and jump in microgravity using a combination of leg motion and thrusters.

### v2

@@ III-E Laboratory Experiment Setup
To validate our approach in a real-world scenario, we conducted experiments using the physical air bearings platform [24] located within the ZeroG Laboratory at the University of Luxembourg. This specific platform floats on an epoxy floor, weighs 5.32 kg and measures 31 cm in radius and 45 cm in height. It is equipped with a Raspberry Pi 4 for onboard control and communication. The ZeroG Lab contains an Optitrack Motion Capture System (MCS) that precisely tracks the platform’s pose at a frequency of 200 Hz. We derive linear and angular velocities through simple forward differencing, that estimate the rate of change of positions and orientations over consecutive time-steps. Thanks to the relatively high accuracy of the MCS, and a reasonable averaging window, concerns about noise sensitivity are negligible. Our experimental setup maintains a connection between a laptop, the MCS, and the FP through a local network. The laptop serves as the ROS (Robot Operating System) master node on the network, subscribing to the Optitrack node to acquire pose data and publishing the actions of the trained agents at a rate of 5 Hz. This action frequency is deliberately constrained to prevent damage to the solenoid valves controlling the thrusters on the floating platform. Figure 3 illustrates the key components interacting during the simulated training and validation phase (on the left), and those interacting during the closed-loop control tests of the real FP system in the Lab (on the right).
@@ IV Experimental Setup
Our experiments encompass both numerical simulation-based evaluations and real-world validations. For the evaluation, each trained policy was tested across a diverse set of scenarios defined by various environmental conditions.
IV-A Performance Metrics
@@
To evaluate the pose task in the lab, we only use the position and orientation error, since we do not have enough experiments to compile more complete statistics. However, we do provide complete trajectories to better understand the behavior of the RL agent and LQR controller.
Finally, for the velocity tracking, we chose to apply the controllers on a trajectory tracking task. For that, we wrote a simple trajectory tracker, that generates a velocity vector to track, based on a sequence of points to follow. This vector is computed by taking the closest point that intersect with a circle of radius $r$ centered around the system. This radius, is a look-ahead-distance which can be tuned to adjust the speed of the tracker. The velocity is considered fixed for the whole of the trajectory, meaning that the instructed velocity is not reduced even if there are sharp corners. This controller is then applied on 3 shapes, a circle, a square and a infinite. For these trajectories, we measure the error in velocity, and the averaged trajectory tracking error.
IV-B Real-World Experimental Validations
To validate the real-world applicability of our simulation-trained control policies, we used the physical floating platform with the laboratory setup described in section III-E to perform a series of experiments. Each test run, for the same policy, initiated the FP from different initial conditions, namely position and orientation within the lab.
@@ V Results
VN UF TD RTF PT5 PT2 PT1 OT5 OT2 OT1 ALV AAV AAS
[…]
Ideal RL - - - - 64 34 6 94 89 73 0.08 0.12 0.29
LQR - - - - 73 41 17 27 11 5 0.07 0.16 0.10
Velocity Noise RL 0.02 - - - 64 30 7 94 90 72 0.08 0.12 0.31
RL 0.04 - - - 61 21 6 94 89 66 0.09 0.13 0.31
LQR 0.02 - - - 53 21 6 4 1 0 0.09 0.49 0.23
LQR 0.04 - - - 14 3 0 2 1 0 0.15 0.56 0.29
Constant Torque RL - - 0.05 - 63 24 2 94 86 61 0.08 0.12 0.35
LQR - - 0.05 - 57 20 6 3 1 0 0.07 0.43 0.35
Constant Force RL - 0.20 - - 63 29 7 94 90 74 0.09 0.12 0.30
RL - 0.40 - - 52 19 5 94 89 72 0.09 0.12 0.31
LQR - 0.20 - - 66 17 4 28 12 6 0.07 0.15 0.12
LQR - 0.40 - - 23 0 0 30 13 6 0.08 0.16 0.15
Constant Force & Torque RL - 0.20 0.05 - 62 24 5 94 86 61 0.08 0.12 0.35
LQR - 0.20 0.05 - 13 2 0 3 1 0 0.07 0.44 0.32
Thruster Failures RL - - - 1 32 15 6 70 55 36 0.10 0.12 0.28
RL - - - 2 15 6 2 45 31 20 0.16 0.15 0.25
LQR - - - 1 40 17 5 20 8 4 0.10 0.21 0.16
LQR - - - 2 12 4 1 11 4 2 0.14 0.28 0.22
V-A Numerical Simulation RL & LQR
In this section, we explore the behaviour of an RL agent trained to perform the “go to pose” task, and compare it to the LQR controller. We chose the “go to pose” task as it is a representative example, allowing us to assess the behaviour of different policies while controlling both the position and the orientation of the FP. To characterize the controllers’ behaviors we expose them to a range of disturbances. Neither the RL agents nor the LQR are specifically adapted to incorporate methods from robust RL or robust optimal control theory. Yet, it is important to acknowledge that the RL agent was trained with some domain randomization to learn how to deal with force disturbances up to 0.25 N. Both of them are evaluated in MuJoCo, with similarly randomized initial conditions. In Table II, each line corresponds to an experiment, with various disturbances applied, and was compiled using 256 trajectories of 250 steps each.
@@ V-A Numerical Simulation RL & LQR
When considering the Velocity Noise (VN), it is observed that with the lowest noise level, the RL performances remain unchanged, while the LQR struggles, in particular with attitude control. With 0.04 m/s of noise, the performance of both controllers decreases. However, the RL controller is more resilient than the LQR controller to this kind of disturbance, even though it was not trained for it. In the interest of brevity, we do not report action noise value in the table, as we found their effect to be negligible on both controllers.
[…]
@@
Overall, while the LQR controller demonstrates greater efficiency and precision in position control with our current tuning, it encounters challenges when subjected to the selected range of disturbances. In contrast, RL exhibits a lower degree of energy conservation but offers stronger resilience when subject to a wide range of disturbances. It is possible that with a different cost function, better tuning of its weights, and a robust optimal control approach, the LQR becomes adept with these disturbances. Similarly, the RL agent could be induced to learn more conservative policy that uses less actions throughout the episodes, via adequate reward shaping. However, the RL agent is not using a robust RL approach either, and domain randomization was only applied on force disturbances up to 0.25N, which is less than the disturbances it can overcome.
@@ V-B ZeroG Laboratory
For experiments with the real FP system, we report tests using both the RL and LQR methods for the “go to pose” task, and tests using the RL agent only for the “track velocity” task.
V-B1 Go to pose
@@
The controllers are run on the FP, which is connected to a constant air supply through a tether. This tether applies some light unknown disturbances such as a small torque and force to the platform. Moreover, the system velocities are derived from the optitrack system. The observed velocities include minor noise and small delays due to network communication.
@@
Figure 4 illustrates the performance of each controller. The first row shows the trajectories of the FP, and the second row shows the distance to the goal in position and orientation. The first two columns have the rough same initial pose: Init1, while the two last share the same initial pose: Init2.
@@
From the last row, it is evident that the LQR controller converges faster in position than the RL controller. This aligns well with the behaviours observed in the simulation benchmark, with an LQR controller converging faster. However, it is also apparent that the LQR solution exhibits a minor overshoot. Such an observation is also in line with the simulation benchmark, as the uneven floor in the lab likely disrupts the LQR controller by applying a subtle constant force, preventing it from reaching its simulation baseline performance. Looking at the top row, we can see that the LQR is also overshooting a bit. Of course, the behaviour can be adjusted by modifying the weights associated with the importance of the error in position in the cost matrix. It is also worth noting that the LQR controller is sensitive to the weights; smaller weights do not incentivize the FP motion toward the goal. In comparison to the simulation, it was deemed necessary to alter the weights of the LQR controller to yield a more aggressive approach to achieve satisfying performances. As for the RL agent, it is noticeable that the FP initially aligns its heading and then gradually converges toward the goal. Consistently with the results from the simulation, the RL controller is significantly more accurate in terms of heading while achieving a position accuracy similar to that of the LQR controller. Overall, both controllers performed well in the lab, reaching their expected performances.
Figure 4: Comparison of the RL and LQR controller on two different initial poses in the ZeroG lab. Init 1 (resp. 2) denotes the first (resp. second) initial pose. For the trajectories, the scale of the y-axis is represented as a log value for better visualization.
Figure 5: RL agent performing velocity tracking in simulated (bottom) and lab (top) environments.
V-B2 Track velocity
In the tests performed for this task in the lab, the objective is to assess the simulation-trained policy ability to adhere to a set of predetermined target velocities. Since the LQR model relies on both position and velocity states as input, while the RL agent only requires velocity, we opted to present the RL policy results for this specific task. Both numerical-simulation and lab tests are displayed to validate the sim-to-real transfer.
Similar to the “go to pose” experiments, the FP was subjected to un-modeled disturbances affecting both linear and angular motion. An additional challenge in these tests was the accurate estimation of velocities, affected by slight measurement noise and communication delays. The pre-generated trajectories to be tracked by the policy were designed to test the FP’s response accuracy and agility.
Figure 5 illustrates the target trajectory and the FP’s position for the circle, square and infinite shapes. It is clearly visible that the hardest task was to follow a squared-shaped trajectory. This is due to the sharp turns that require precise maneuvering and acceleration adjustments, which could be induced by reducing the look-ahead-distance and target velocity of the tracking when close to corners. The performance metric used is the linear velocity error $e_{v}$ expressed as $\mu\pm\sigma$, where $\mu$ is the mean and $\sigma$ is the standard deviation during the test duration. Table III reveals that the lab environment generally presents higher velocity errors compared to the simulation environment, particularly notable in the square shape with a lab error of 0.07 $\pm$ 0.05 m/s versus a sim error of 0.05 $\pm$ 0.08 m/s, the difficulty of real-world transfer. For the infinite trajectory, we observed a slight overshoot in the path’s lower regions, caused by the irregularities in the epoxy floor, which are significant in that area of the laboratory, affecting the FP’s motion. This can also be seen on the square, and to less of a degree on the circle. In our case, there is a slope pulling free-floating objects towards negative y.
Shape Lab Error ($\mu\pm\sigma$) [m/s] Sim Error ($\mu\pm\sigma$) [m/s]
circle 0.03 $\pm$ 0.02 0.01 $\pm$ 0.01
infinite 0.04 $\pm$ 0.03 0.01 $\pm$ 0.01
square 0.07 $\pm$ 0.05 0.05 $\pm$ 0.08
TABLE III: Comparison of Velocity Errors Between Lab and Sim Environments for the track velocity task. All the trajectories are tracked at 0.2 m/s.
@@ VI Conclusions
This study presents a robust framework for the 2D control of a FP using deep reinforcement learning techniques. The enhancements to our simulator extend its applicability, enabling the training of agents for complex tasks. The demonstrated effectiveness of the PPO algorithm, with seamless transfer behavior from the simulation to the FP system, highlights its potential for autonomous navigation in space. Furthermore, our comparative analysis with traditional optimal control algorithms reveals the superior performance of our method in the presence of unpredictable stochastic disturbances. In the future, we aim to explore the efficacy of complex architectures such as LSTMs or Transformers to see if they can cope better with a greater number of disturbances and in particular actuators failure. Additionally, we aim to broaden our framework to train agents capable of walking and jumping in microgravity using a combination of leg motion and thrusters, expanding the set of available sensors and actuators.
APPENDIX
VI-A PPO Hyperparameters
Table IV outlines the key parameters used in the adapted version of the Proximal Policy Optimization (PPO) algorithm for training our models.
Parameter Value
Algorithm PPO
Network Type Actor-Critic MLP
Separate Networks True
MLP Units [128, 128]
Activation Function tanh
Initializer Identity
Regularizer None
Learning Rate $41e-$ 4
Gamma ($\gamma$) 0.99
Tau ($\tau$) 0.95
Entropy Coefficient 0.0
Horizon Length 16
Minibatch Size 8192
Mini Epochs 8
Critic Coefficient 0.5
Gradient Clipping Norm 1.0
KL Threshold 0.016
Critic Coefficient 0.5
TABLE IV: PPO Training Parameters
VI-B LQR parameters
Table V summarizes the parameters of the Discrete LQR Controller used. The controller is made planar compatible, indicating a restriction to the 2D plane.
Parameter Value
Name $LQR$
Q (State cost matrix) [0.0001, 1e-05, 100, 100, 1e-06, 1e-06, 1]
R (Control cost matrix) [0.01, 0.01, 0.01, 0.01, 0.01, 0.01, 0.01, 0.01]
W (Disturbance weight matrix) [0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1]
Make planar compatible Yes
Control type LQR
TABLE V: Parameters for the Discrete LQR Controller
ACKNOWLEDGMENT
Work supported by the European Union’s Horizon 2020 research and innovation program under grant No. 101096487.